# G2SDF: Surface Reconstruction from Explicit Gaussians with Implicit SDFs

### v1

@@ Abstract
State-of-the-art novel view synthesis methods such as 3D Gaussian Splatting (3DGS) achieve remarkable visual quality. While 3DGS and its variants can be rendered efficiently using rasterization, many tasks require access to the underlying 3D surface, which remains challenging to extract due to the sparse and explicit nature of this representation. In this paper, we introduce G2SDF, a novel approach that addresses this limitation by integrating a neural implicit Signed Distance Field (SDF) into the Gaussian Splatting framework. Our method links the opacity values of Gaussians with their distances to the surface, ensuring a closer alignment of Gaussians with the scene surface. To extend this approach to unbounded scenes at varying scales, we propose a normalization function that maps any range to a fixed interval. To further enhance reconstruction quality, we leverage an off-the-shelf depth estimator as pseudo ground truth during Gaussian Splatting optimization. By establishing a differentiable connection between the explicit Gaussians and the implicit SDF, our approach enables high-quality surface reconstruction and rendering. Experimental results on several real-world datasets demonstrate that G2SDF achieves superior reconstruction quality than prior works while maintaining the efficiency of 3DGS.
@@ 1 Introduction
Figure 1: G2SDF. Our method establishes a connection between the implicit SDFs and explicit Gaussians, linking the opacity of each Gaussian to its distance from the surface. Designed to handle unbounded scenes, our approach achieves superior surface reconstruction, particularly in areas with limited view coverage. Compared to 2DGS [1] and GOF [2], our method achieves higher F1 scores ($\uparrow$) and reconstructs smooth surfaces with fine details.
@@
3D Gaussian Splatting (3DGS) [3] has emerged as a state-of-the-art method for high-quality novel view synthesis by leveraging the rasterization pipeline and representing 3D scenes using points characterized by Gaussian functions. However, 3DGS remains an explicit and discrete representation, posing challenges for accurate 3D surface reconstruction. This limitation is in particular important for applications such as geometry editing [4, 5, 6], 3D animation [7, 8, 9], and robotics [10, 11, 12]. The primary challenge lies in 3DGS being a discrete and unstructured point-based geometry representation. Given that the optimization process is only guided by the rendering quality, the Gaussian representation tends to exhibit noisy geometry of the scene which complicates the extraction of accurate 3D surfaces through post-processing methods like Poisson surface reconstruction [13].
These challenges have motivated recent investigations to explore how 3DGS can be used for high-quality surface reconstruction while maintaining its rendering speed and efficiency. 2DGS [1] employs 2D instead of 3D Gaussians, forcing Gaussians to align closer to surface. They extract meshes through Truncated Signed Distance Field (TSDF) fusion [14]. However, TSDF fusion struggles with accurately modeling thin structures and reconstructing unbounded scenes, leading to rendered depth maps which can be inaccurate. SuGaR [15] addresses these issues by coarse-to-fine reconstruction. It regulates 3D Gaussians to align with coarse mesh faces, then uses Poisson surface reconstruction [13] in second step to extract a mesh, although Poisson reconstruction ignores the Gaussians’ opacities and scales. Overall, the primary challenge remains in the inconsistency between mesh extraction and Gaussian rasterization during training. Gaussian Opacity Fields (GOF) [2] attempts to address this by generating an opacity field from Gaussians and extracting surfaces using Marching Tetrahedra [16]. Nevertheless, the opacity field is created via post-processing through ray tracing, which requires a specific Gaussian rasterizer. This approach demands extra time and resources, and still suffers from inconsistency between the underlying geometry and Gaussians.
To achieve high-quality reconstructions while preserving the efficency of 3DGS, we introduce G2SDF, a novel approach that seamlessly integrates an implicit SDF into the Gaussian Splatting optimization. Our contributions are:
  • A hash-based SDF which provides spatial continuity to the originally unstructured Gaussians without additional volume rendering training. Rather than solely optimizing Gaussian positions through appearance supervision, we use the SDF to effectively constrain their spatial distribution, creating a more cohesive 3D representation.
  • A differentiable SDF-to-Opacity function that links the implicit SDF with Gaussian opacities, ensuring Gaussians closer to the surface exhibit higher opacity values and enforcing the distribution of Gaussians to be more surface aligned.
  • A novel coordinate normalization function for our hybrid representation that maps unbounded 3D coordinates to a 0-1 range, enabling efficient positional embedding.
We find that the joint hybrid optimization of an implicit SDF and the explicit Gaussian Splatting representation leads to better reconstructions while maintaining 3DGS’s efficiency. Experiments on real-world unbounded datasets demonstrate that our method delivers high-quality reconstructions while preserving the view synthesis quality. Note that our method is also easily adaptable to any existing Gaussian rasterizer and can be used with different methods for mesh extraction such as Marching Cubes [17], Marching Tetrahedra [16] and TSDF fusion [14].
Figure 2: Overview. Our hybrid representation consists of the Gaussian Splatting pipeline and an SDF network, integrating implicit SDFs into the explicit Gaussian representation. Our SDF network first normalizes any 3D input to (0,1) range and predicts an SDF value. Then using a differentiable SDF-to-Opacity function, we connect the SDF value $\bar{s}$ with the opacity value $\alpha$ and force Gaussians to follow a certain distribution near the surface. Rendered depth is used as a pseudo-ground truth signal to supervise the SDF directly in 3D. We further apply an off-the-shelf depth estimator to supervise rendered depth and normal.
@@ 2.1 View Synthesis and Gaussian Splatting
Neural Radiance Fields (NeRF) [18] utilize multi-layer perceptions (MLP) as scene representation to predict geometry and view-dependent appearance. The MLP is optimized via a photometric loss through volume rendering. Subsequent methods have focused on optimizing NeRF’s training and expressiveness using grid representations [19], improving rendering speed [20, 21] and scaling to unbounded scenes [22]. However, volume rendering typically requires substantial computational resources and extensive training durations. Recently, 3D Gaussian Splatting (3DGS) [3] has emerged as an efficient alternative for view synthesis. By leveraging the differentiability of Gaussian functions, 3DGS achieves high-quality rendering while enabling real-time and efficient novel view image synthesis. Subsequent works have enhanced rendering quality through anti-aliasing techniques [23, 24, 25] or improved rendering speed by either controlling the density of Gaussians or using radiance field as a prior [26, 27, 28]. DNGaussian [29] introduces a depth-regularized approach to address the challenge of geometry degradation in sparse input views. GeoGaussian [30] focuses on preserving the geometry of non-textured regions like walls and furniture, which tend to degrade over time. Instantsplat [31] uses Dust3r [32] to initialize the point cloud and achieves fast training on sparse view case. Instead of representing scenes with explicit Gaussians, Scaffold-GS [33] utilizes a combination of MLPs and feature grids to infer Gaussians for rasterization, merging implicit and explicit representations with a focus on view synthesis. However, these methods primarily focus on enhancing appearance quality while neglecting underlying geometry, limiting their application to view synthesis only.
@@ 2.2 Surface Reconstruction from Gaussians
Due to the discrete and unstructured nature of 3DGS, along with supervision being limited to RGB images during training, previous methods often struggle to accurately capture scene geometry, making surface reconstruction a challenging task. SuGaR [15] addresses this by building a density field from Gaussians aggregating the opacity values of nearby Gaussians for any given 3D point. It extracts meshes from this density field using level-set searching and employs a two-stage training process where meshes from the coarse stage are used to guide the alignment of Gaussians. However, SuGaR is time-consuming and struggles with reconstructing large, smooth areas such as floors or walls. Other methods [34, 35, 36, 37, 38] employ either depth or normal estimators as priors to supervise Gaussian optimization. They then use TSDF fusion or Poisson reconstruction to extract meshes from multi-view depth maps or Gaussians. However, TSDF fusion faces challenges in handling unbounded scenes, whereas Poisson reconstruction ignores all Gaussian attributes. 2DGS [1] partly addresses this by flattening 3D Gaussians into 2D disks which better align with the underlying surfaces, improving geometric accuracy. Gaussian Opacity Fields (GOF) [2] constructs an opacity field by approximating the minimum accumulated alpha values from all views using ray tracing and employs a depth distortion loss from 2DGS [1] to improve smooth surface reconstruction. Surface extraction is done using Marching Tetrahedra [16], but the distribution of Gaussians relative to scene surfaces remains inconsistent and unstructured. 3DGSR [39] improves upon this by linking each Gaussian’s opacity value with its distance to the surface through a hash-grid based SDF. However, to effectively train the SDF, it incorporates an additional branch to volume render additional depth and normal maps supervised by depth and normal from Gaussian rasterization, which is in-efficient and the inconsistent rendering may degrade both branches. Our hybrid representation integrates a hash-grid into the 3DGS representation and is trained jointly, preserving both the speed and quality of training and rendering, while enabling detailed reconstruction.
@@ 3 Method
3.1 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [3] employs a set of 3D points to effectively render images from given viewpoints, each characterized by a Gaussian function with 3D mean $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, opacity value $\alpha_{i}\in\mathbb{R}$, RGB color values $\mathbf{c}_{i}\in\mathbb{R}^{3}$:
@@
where $T_{i}=\prod_{j=1}^{i-1}(1-o_{j})$, depth $d_{i}$ is the distance between camera center and the ray-Gaussian intersection plane, and Gaussian’s normal $\mathbf{n}_{i}$ is approximated the as the normal of the ray-Gaussian intersection plane.
3.2 Neural Signed Distance Field
In our hybrid representation, we optimize a neural signed distance field (SDF) next to the 3DGS representation. To scale to unbounded scenes and achieve high fidelity, we combine multi-layer perceptrons (MLPs) with grid-based backbones [41] as outlined in the following.
Unbounded Normalization. Grid-based representations require defining a specific scene bounding box to initialize the grid’s volume size. To adapt to unbounded real-world scenes at various scales, we first apply a sigmoid-like normalization function on any 3D point $\mathbf{x}\in\mathbb{R}^{3}$, mapping any scope to a fixed $(0,1)$ range:
$h(\mathbf{x})=1/(1+exp(-\sigma\mathbf{x})),$ (4)
where $\sigma=2/B$, $B\in\mathbb{R}^{3}$ is the bounding box size of initial Gaussians. This term ensures that the volume containing the majority of Gaussians occupies most of the grid. During Gaussian densification, some new Gaussians may extend beyond this primary bounding box but remain within the defined scope $(0,1)$ after normalization.
Parameterization. To better encode the scene geometry, we choose One-blob $\gamma(\mathbf{x})$ [42] as positional embedding and a multi-resolution hash-based feature grid $\mathcal{V}=\{\mathcal{V}^{l}\}^{L}_{l=1}$ [19]. Feature $\mathcal{V}(h(\mathbf{x}))$ at any 3D point $\mathbf{x}$ are queried via trilinear interpolation. The geometry decoder $f$ is a MLP which maps the 3D coordinate to a SDF value $\bar{s}$:
@@ 3.2 Neural Signed Distance Field
Therefore, for each 3D Gaussian located at $\mathbf{\mu}_{i}$, its corresponding SDF value can be written as $\bar{s}_{i}=f(\mathbf{\mu}_{i})$.
3.3 Scene Representation and Optimization
While 3DGS is an explicit and discrete representation, SDFs provide an implicit representation that facilitates precise surface extraction. In this section, we discuss how to leverage the advantages of both and optimize them jointly.
Hybrid Scene Representation. We propose a hybrid scene representation that combines the strengths of neural implicit SDF with 3DGS, as shown in Fig. 2. First, an SDF value is predicted for each Gaussian using Eq. 5, and it is then converted to opacity value which is needed by the Gaussian rasterizer. We use a differentiable Gaussian-shaped transformation $g(\bar{s})$:
$g(\bar{s}_{i})=\exp(-\beta*\bar{s}_{i}^{2})\mapsto\alpha_{i},$ (6)
where $\beta$ is a learnable scale factor. This transformation has two main properties that ensures Gaussians to follow a structured distribution. First, it establishes a connection between the opacity and the position of a Gaussian Second, it enforces Gaussians to remain close to the surface, as their opacity values drop to zero with increasing distance.
Joint Optimization. To train such a hybrid model, a straightforward approach is to enforce all Gaussians to meet the condition $\bar{s}_{i}=0$, resulting in $\alpha_{i}=1$. But experimentally we found that this negatively impacts rendering quality as achieving high-quality renderings requires variation in Gaussian opacity and providing SDF supervision solely at the position of Gaussians is too sparse. On the other hand, supervising the SDF through volume rendering like 3DGSR [39] introduces substantial computational overhead, with supervision signals limited to 2D, which is insufficient for accurate SDF representation.

### v2

# MonoGSDF: Exploring Monocular Geometric Cues for Gaussian Splatting-Guided Implicit Surface Reconstruction

@@ Abstract
Accurate meshing from monocular images remains a key challenge in 3D vision. While state-of-the-art 3D Gaussian Splatting (3DGS) methods excel at synthesizing photorealistic novel views through rasterization-based rendering, their reliance on sparse, explicit primitives severely limits their ability to recover watertight and topologically consistent 3D surfaces. We introduce MonoGSDF, a novel method that couples Gaussian-based primitives with a neural Signed Distance Field (SDF) for high-quality reconstruction. During training, the SDF guides Gaussians’ spatial distribution, while at inference, Gaussians serve as priors to reconstruct surfaces, eliminating the need for memory-intensive Marching Cubes. To handle arbitrary-scale scenes, we propose a scaling strategy for robust generalization. A multi-resolution training scheme further refines details and monocular geometric cues from off-the-shelf estimators enhance reconstruction quality. Experiments on real-world datasets show MonoGSDF outperforms prior methods while maintaining efficiency.
@@ 1 Introduction
Figure 1: MonoGSDF. Ours achieves superior surface reconstruction on arbitrary-scale scenes, particularly in areas with limited view coverage. Compared to 2DGS [12] and GOF [48], ours achieves higher F1 scores ($\uparrow$) and reconstructs smooth surfaces with fine details.
@@
3D Gaussian Splatting (3DGS) [16] has emerged as a state-of-the-art method for high-quality novel view synthesis by leveraging the rasterization pipeline and representing 3D scenes using points characterized by Gaussian functions. However, 3DGS remains an explicit and discrete representation, posing challenges for accurate 3D surface reconstruction and meshing. This limitation is particularly important for applications such as geometry editing [43, 6, 36], 3D animation [33, 49, 30], and robotics [56, 15, 50]. The primary challenge lies in 3DGS being a discrete and unstructured point-based geometry representation. Given that 3D Gaussians are primarily optimized for accurate appearance rendering rather than geometric regularization, which complicates the extraction of accurate 3D surfaces through post-processing methods like Poisson surface reconstruction [14].
These challenges have motivated recent investigations exploring how 3DGS can be used for high-quality surface reconstruction while maintaining its rendering speed and efficiency. Methods like 2DGS [12] address this by aligning 2D Gaussians to surfaces, although their depth map-based mesh extraction [7] struggles with thin structures and unbounded scenes due to resolution and resource limitations. Methods such as SuGaR [11] attempt to address these limitations by aligning 3D Gaussians with mesh faces, while DN-Splater [34] incorporates depth and normal estimators. Nevertheless their reliance on Poisson reconstruction [14] still fundamentally limits reconstruction and meshing quality due to the inherent geometric inaccuracies of Gaussian representations. Although recent methods [45, 24, 55] have integrated SDF fields into Gaussian training for surface reconstruction, their incomplete exploration of the Gaussian-SDF relationship results in over-smoothed surfaces and limits their application within object-level reconstruction.
To bridge the gap between explicit Gaussian representations and implicit surface reconstruction while fully leveraging their complementary advantages, we present MonoGSDF, a unified framework that establishes a robust connection between 3D Gaussians and SDF-based surface modeling for high-fidelity reconstruction. Our contributions are:
  • A Gaussian-based implicit surface reconstruction pipeline for monocular images, where the SDF guides Gaussians to better align near the surface during training, while Gaussians act as priors for fast zero-level set extraction during reconstruction.
  • A simple yet efficient SDF-to-Opacity function that bridges implicit SDF with explicit Gaussians, coupled with a Gaussian-guided normalization strategy that dynamically maps unbounded 3D coordinates to a (0,1) range with minimal distortion.
  • A multi-resolution training strategy based on wavelet transforms progressively refines details, while off-the-shelf geometric cues further enhance reconstruction quality, enabling Gaussians to better capture scene geometry.
Experiments on real-world datasets demonstrate that our method delivers high-quality reconstructions and outperforms baselines by 13% in terms of Chamfer distance while preserving the view synthesis quality.
Figure 2: Overview. MonoGSDF establishes a synergistic relationship between SDF and Gaussian representation: the SDF regulates Gaussian distribution during training, while Gaussians serve as efficient primitives for rapid surface extraction. We first query the SDF values for each Gaussians and convert them into opacity values. This enforces Gaussian to stick closer to the surface. Then we follow the standard rasterization pipeline to render color, depth and normal maps. Off-the-shelf geometry cues and multi-resolution strategy are applied for better geometry performance. And we jointly train the SDF via 3D supervision signal from rendered Gaussian depth.
@@ 2.1 View Synthesis and Gaussian Splatting
Neural Radiance Field (NeRF) [25] utilize multi-layer perceptions (MLP) as scene representation to predict geometry and view-dependent appearance. The MLP is optimized via a photometric loss through volume rendering. Subsequent methods have focused on optimizing NeRF’s training and expressiveness using grid representations [28], improving rendering speed [32, 53] and scaling to unbounded scenes [3]. However, volume rendering typically requires substantial computational resources and extensive training durations. 3D Gaussian Splatting (3DGS) [16] has emerged as an efficient approach for real-time view synthesis through differentiable Gaussian functions. Subsequent works have enhanced rendering quality via anti-aliasing techniques [47, 40, 31] and improved rendering speed through Gaussian density control [42] and radiance field priors [22, 29]. Geometry-focused approaches like DNGaussian [19] address sparse view degradation, while GeoGaussian [20] preserves non-textured regions. Instantsplat [9] accelerates sparse view training using Dust3r [38] initialization, and Scaffold-GS [23] combines implicit-explicit representations. However, these methods primarily focus on appearance quality rather than underlying geometry, limiting their application to view synthesis.
@@ 2.2 Surface Reconstruction from Gaussians
Due to the discrete and unstructured nature of 3DGS, along with supervision being limited to RGB images during training, existing methods often struggle to accurately capture scene geometry, making surface reconstruction particularly challenging. SuGaR [11] addresses this by constructing a density field from Gaussians and extracting meshes via level-set searching. However, it is computationally expensive and struggles to reconstruct large, smooth surfaces like floors and walls. Other approaches [4, 5, 51, 34, 39] incorporate depth or normal estimators as priors to supervise Gaussians. 2DGS [12] and GSurfels [8] improves geometric alignment by flattening 3D Gaussians into 2D disks or surfels, enabling better surface representation. For surface reconstruction, they rely on either multi-view depth maps for TSDF fusion or Poisson reconstruction. However, TSDF fusion is constrained by fixed resolution and high memory usage for large scenes. To handle unbounded scenes, it requires contracting coordinates, introducing unnecessary distortions. Meanwhile, Poisson reconstruction often produces noisy surfaces due to inaccuracies in Gaussian geometry distribution. Instead of refining depth and normal maps, some methods [45, 55, 24, 2] integrate Signed Distance Field (SDF). 3DGSR [24] and GSDF [45] associate Gaussians with SDF via an additional branch that volume-renders depth and normal supervised by rendered Gaussian depth and normal. However, this approach is inefficient and fails to fully exploit the relationship between Gaussians and SDF. GS-Pull [55] and GSurf [2] improve upon this by leveraging SDF gradients to better align Gaussians with surfaces, but they are either restricted to object-level reconstruction or produce overly smooth results. Gaussian Opacity Fields (GOF) [48] approximates minimum accumulated alpha values from all views via ray tracing to construct an opacity field, followed by surface extraction using Marching Tetrahedra [18]. However, Gaussian distributions remain inconsistent and unstructured relative to scene surfaces. Our method leverages SDF to guide Gaussians to align closer to the surface during training. At inference, Gaussians serve as primitives to efficiently constrain zero-level set extraction, avoiding redundant free-space searches while achieving accurate reconstruction.
@@ 3 Method
3.1 Preliminaries
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) [16] employs a set of 3D points to effectively render images from given viewpoints, each characterized by a Gaussian function with 3D mean $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, opacity value $\alpha_{i}\in\mathbb{R}$, RGB color values $\mathbf{c}_{i}\in\mathbb{R}^{3}$:
@@ 3.1 Preliminaries
where $T_{i}=\prod_{j=1}^{i-1}(1-o_{j})$, depth $d_{i}$ is the distance between camera center and the ray-Gaussian intersection plane, and Gaussian’s normal $\mathbf{n}_{i}$ is approximated as the normal of the ray-Gaussian intersection plane.
Neural Signed Distance Field. While 3DGS excels in novel view synthesis, its discrete nature limits surface reconstruction. In contrast, SDF provide a continuous, watertight surface as the zero-level set. We use a hash-based neural representation and jointly train the SDF with 3DGS, enabling more precise and efficient surface extraction. To better encode the scene geometry, we choose One-blob $\gamma(\mathbf{x})$ [26] as positional embedding and a multi-resolution hash-based feature grid $\mathcal{V}=\{\mathcal{V}^{l}\}^{L}_{l=1}$ [28]. Feature $\mathcal{V}(h(\mathbf{x}))$ at any 3D point $\mathbf{x}$ are queried via trilinear interpolation. The geometry decoder $f$ is an MLP which maps the 3D coordinate to an SDF value $\bar{s}$:
@@
where $h$ is a coordinate normalization. Therefore, for each 3D Gaussian located at $\mathbf{\mu}_{i}$, its corresponding SDF value can be written as $\bar{s}_{i}=f(\mathbf{\mu}_{i})$.
3.2 Gaussian-Guided Signed Distance Field
Connect Gaussians with SDF. As illustrated in Figure 2, our method begins by querying SDF values with each Gaussian 3D mean $\mathbf{\mu}_{i}$ using Eq. 4, where zero values correspond to surface points. While 3DGSR [24] introduced a bell-shaped function to map SDF value to Gaussian opacity, this approach assigns an opacity value $\alpha_{i}=0.25$ to surface points ($\bar{s}=0$), which contradicts the physical reality that surface points should exhibit maximum opacity ($\alpha_{i}=1$). To address this limitation, we propose a Gaussian-shaped transformation that maintains both mathematical simplicity and computational efficiency while ensuring physically plausible opacity values $g(\bar{s})$:
$g(\bar{s}_{i})=\exp(-(\beta*\bar{s}_{i})^{2})\mapsto\alpha_{i},$ (5)
where $\beta$ is a hyperparameter. While previous 3DGS methods [16, 12, 23] optimize opacities through alpha blending without considering spatial distribution constraints, these approaches often result in limited geometric accuracy. Our method enforces a better distribution of Gaussians along the surface, leading to high geometric fidelity.
Gaussian-Guided Normalization Our rendering pipeline follows the standard Gaussian rasterization process, generating color $\mathbf{\bar{C}}$, depth $\bar{D}$ and $\mathbf{\bar{N}}$ maps from Gaussian attributes, with the depth map subsequently used for SDF training. While 3DGSR [24] and GSDF [45] have adopted similar approaches, their grid-based representations necessitate predefined scene bounding boxes for volume initialization, restricting their applicability to object-level or small-scale scene reconstruction. Although MipNeRF [3] introduced a fixed coordinate contraction method for unbounded scenes, its mapping of infinite space to [-2, 2] conflicts with our hash-based feature field’s requirement for (0, 1) input. Moreover, MipNeRF’s linear transformation within a limited region of interest introduces substantial distortion in peripheral areas. To address these limitations and to enable scalable reconstruction of real-world unbounded scenes, our method introduces a Gaussian-guided normalization strategy. We first establish a bounding box based on the Gaussians, then apply a novel sigmoid-like normalization function that maps 3D point $\mathbf{x}\in\mathbb{R}^{3}$ to a fixed $(0,1)$ range:
$h(\mathbf{x})=1/(1+\exp(-\sigma\mathbf{x})),$ (6)
where $\sigma=2/B$, $B\in\mathbb{R}^{3}$ is the bounding box size of initial Gaussians. This approach ensures the volume containing the primary Gaussians occupies the majority of the grid space. During Gaussian densification, while new Gaussians may extend beyond the initial bounding box, they remain within the normalized scope. Our sigmoid-like normalization maintains near-linear transformation for Gaussians within the primary bounding box while effectively contracting distant Gaussians, achieving minimal distortion.
3.3 Optimization
Training our Gaussian-guided SDF presents unique challenges: enforcing all Gaussians to satisfy the condition $\bar{s}_{i}=0$ degrades rendering quality by removing opacity variations, while sparse Gaussian-based SDF supervision lacks spatial coverage. Volume rendering methods like 3DGSR [24] introduce computational overhead and limited 2D supervision is insufficient for accurate 3D representation. Our solution introduces a novel framework that enhance the Gaussians with an off-the-shelf geometry estimator and supervise the SDF with a dense 3D regularization through depth back-projection.
Geometry Cues. To enhance reconstruction quality in real-world scenarios, we integrate an off-the-shelf monocular depth estimator [^1] We use DepthAnything V2 [41] in this work. to obtain dense per-pixel disparity priors as $\hat{Z}$. However, these estimated pseudo disparity measurements exhibit significant scale ambiguity and suffer from multi-view inconsistency. While DN Splatter [34] addresses this issue through direct depth alignment, applying a transformation $\bar{D}=s\hat{D}+t$, their method fails to account for shifts in disparity $\hat{Z}$, resulting in distance-dependent scale variations. To ensure robust relative scale consistency, we first utilize the rendered depth to correct the pseudo disparity and obtain the pseudo depth $\hat{D}=a/(s*\hat{Z}+t)+b$:
$\mathcal{L}_{D}=\sum|({a/(s*\hat{Z}+t)+b-\bar{D}}|.$ (7)
where $s,t$ and $a,b$ represent learnable scale and shift parameters. We further convert the aligned depth into pseudo normal $\mathbf{\hat{N}}=\nabla\hat{D}/||\nabla\hat{D}||$, providing a more stable supervision signal for Gaussian training:
$\mathcal{L}_{N}=\sum(1-\mathbf{\hat{N}}\cdot\mathbf{\bar{N}}).$ (8)
where $\mathbf{\bar{N}}$ is rendered Gaussian normal. Therefore, the overall geometry regularization is $\mathcal{L}_{geo}=\lambda_{D}\mathcal{L}_{D}+\lambda_{N}\mathcal{L}_{N}$.
While the geometry cues offer valuable 2D supervision for improving depth and normal, they fail to constrain the 3D Gaussian distribution effectively. We address this limitation through the SDF, which enforces tied surface alignment of Gaussians.
Multi-View Consistent 3D Regularization. Directly supervising SDF with monocular geometry cues like MonoSDF [46] suffers from scale ambiguity and lacks multi-view consistency. Further, volume rendering-based supervision as proposed in 3DGSR [24] and GSDF [45] are inefficient and only supervise SDF with rendered Gaussian depth map in 2D. Our approach combines monocular cues with the inherent advantages of 3D Gaussians that the Gaussian-rendered depth $\bar{D}$ naturally maintains multi-view consistency and enables reliable 3D signal supervision for SDF through back-projection.